The sum of the ages of a brother and sister is 27. If four times the brother's age is subtracted from three times the sister's age, the difference is 11.

Answer by ewatrrr(24785)


Hi, The sum of the ages of a brother and sister is 27
Let x and (27-x))represent the sister's and brother's ages respectively
Question states***
3x - 4(27-x)= 11
Solving for x
3x + 4x - 108 = 11
7x = 119
x = 17, the sister's age. The brother is 10 (27-17)

CHECKING our Answer***
3*17 - 4*10 = 51 - 40 = 11